Question 3 1 pts de =k(Q-70) Q = -70 + Cekt O Q = 70 + Cekt Q = -70+ Cek O Q = 70+ Cekt Next Previous

Note: Give me a detailed or step by step solution. Thank you
Answer #1

2 do at K (④-7) ㅋ -KQ-70K di | мкрат dd - kg = -70k at 을This is first order linear non differential equation in homogence 04 the formdy da + Ay=B where A and B are functions of a (or constants). The solution of such differential equation is given by ivegratiAnd solution is given = JB (IF) da te where is constant of integration, GrIF)So, for de -KQ = -70K atofert) = -70kfetto Q (IF) = Se 70k) (IF)dt+ grekt) -- 70K $(TF) 46 +C dttc grekt) =-70K.ekt tc (*) . Year due = eante alekt) = 70ekt How dividing by -kt acekty = 7ēkt = + c ékt ekt - Q Q = t C 70 ēkt ekt éktkt Q = 70+ cel [ 1 =xh x 8= 70+ cektThis matches with and option So, and option is correct Q= 70+ cekt choice. D